Sharon's mom is eight more than twice her age. Her father is 3 years older than her mother. If the three family members have lived a total of 94 years, how old are each family member?
Found 2 solutions by stanbon, mananth:
Answer by stanbon(75887)

Sharon's mom is eight more than twice her age.
---
m = 2s+8
----
Her father is 3 years older than her mother.
---
f = m+3
----
If the three family members have lived a total of 94 years, how old are each family member?
s + m + f = 94
---
Substitute for "s" and for "f" and solve for "m":
(m-8)/2 + m + m+3 = 94
----
m-8 + 2m + 2m+6 = 188
5m = 190
m = 38 (mother's age)
----
f = m+3 = 41 (father's age)
---
s = (m-8)/2 = 15 (sharon's age)
==============
Cheers,
Stan H.

Answer by mananth(16946)

Sharon's mom is eight more than twice her age. Her father is 3 years older than her mother. If the three family members have lived a total of 94 years, how old are each family member?
..
Let Sharon's age be x
Mom = 2x+8
father = 2x+8+3= 2x+11
...
x+2x+8+2x+11=94
5x+19=94
-19 -19
5x=75
/5
x= 15 sharon's age.
Mother = 2x+8= 2*15+8=38 years
Father = 38+3 = 41 years

m.ananth@hotmail.ca